Question 1
YHOO MSFT GOOG BOA WMT NFXL MGM Total
Total Ret 28.576% 42.797% 100.017% 49.643% 77.830% 150.000% 80.000%
$ Investment    100,000       200,000    300,000      50,000    500,000    500,000                200,000
Beta         0.911            0.921         0.904         0.934         0.488         1.350                    1.600
What is the weight for Yahoo in your portfolio?
No short-sales is allowed
Question 2
YHOO MSFT GOOG BOA WMT NFXL MGM Total
Total Ret 28.576% 42.797% 100.017% 49.643% 77.830% 150.000% 80.000%
$ Investment    100,000       200,000    300,000      50,000    500,000    500,000                200,000
Beta         0.911            0.921         0.904         0.934         0.488         1.350                    1.600
What is the return for your portfolio?
Question 3
YHOO MSFT GOOG BOA WMT NFXL MGM Total
Total Ret 28.576% 42.797% 100.017% 49.643% 77.830% 150.000% 80.000%
$ Investment    100,000       200,000    300,000      50,000    500,000    500,000                200,000
Beta         0.911            0.921         0.904         0.934         0.488         1.350                    1.600
What is the beta for your portfolio?
Question 4
YHOO MGM
Total Ret 28.576% 80.000%
$ Investment
Beta         0.911            1.600
No short-sales is allowed
If you manage $10,000,000, how much will you invest in YHOO and in MGM to make the portfolio
Return equal to 50%
Report in Sakai portfolio beta.
(Hint, using the initial value 1 and 1 investment in each stock, calculate Return and Beta and then use solver)
Question 5
YHOO WMT NFXL MGM Total
Total Ret 28.576% 77.830% 150.000% 80.000%
$ Investment                 1                    1                 1                 1
Beta         0.911            0.488         1.350         1.600
No short-sales is allowed
If you manage $10,000,000, find the highest return you can obtain if you want to maintain the
systematic risk of your portfolio equal to 1.00
(Hint, using the initial value 1 and 1 investment in each stock, calculate Return, beta (if needed) and the use solver)
Question 6
YHOO WMT NFXL MGM Total
Total Ret 28.576% 77.830% 150.000% 80.000%
$ Investment                 1                    1                 1                 1
Beta         0.911            0.488         1.350         1.600
No short-sales is allowed
If you manage $10,000,000, find the Lowest systematic risk you can obtain if you want to maintain the
return of your portfolio equal to 85.00%
(Hint, using the initial value 1 and 1 investment in each stock, calculate Return, beta (if needed) and the use solver)
Question 7
MSFT GOOG BOA WMT Total
Total Ret 42.797% 100.017% 49.643% 77.830%
$ Investment                 1                    1                 1                 1
Beta         0.921            0.904         0.934         0.488
No short-sales is allowed
If you manage $10,000,000, find the highest return you can obtain if you want to maintain the
systematic risk of your portfolio equal to 0.6
(Hint, using the initial value 1 and 1 investment in each stock, calculate Return, beta (if needed) and the use solver)
Question 8
MSFT GOOG BOA WMT Total
Total Ret 42.797% 100.017% 49.643% 77.830%
$ Investment                 1                    1                 1                 1
Beta         0.921            0.904         0.934         0.488
No short-sales is allowed
If you manage $10,000,000, find the Lowest systematic risk you can obtain if you want to maintain the
return of your portfolio equal to 85.00%
(Hint, using the initial value 1 and 1 investment in each stock, calculate Return, beta (if needed) and the use solver)
Answer #1

Ans.(1) From details provided in Q.

Weight of Particular Stock= (Amount invested in that stock/ Total Investment in a Portfolio)

So, Weight of YHOO= (100000/1850000)= 0.0541 i.e. 5.41%

Ans.(2) Return from Portfolio=(R1*P1+R2*P2+R3*P3+........+Rn*Pn) i.e. Rate of Return * Investment in Particular Stock

Required Return=(100000*28.58%+200000*42.80%+300000*100.02%+50000*49.64%+500000*77.83%+500000*150%+200000*80%)

=(28,576.00 +85,594.00 +3,00,051.00 +24,821.50 +3,89,150.00 +7,50,000.00 +1,60,000.00 )=$1738192.50 (Ans)

Ans.(3). Beta of Portfolio.=Sum of Weighted Beta of Each Stock i.e. Beta of Particular Stock * Weight of that Stock

Weight of Stock =(Investment in Particular Stock /Total Investment in Portfolio)

Particular YHOO MSFT GOOG BOA WMT NFXL MGM Total

Weight(Investment/Total. Portfolio Value)

0.054054054

  

0.108108108

0.162162162

0.027027

0.27027027

0.27027027

0.108108108

Beta 0.911 0.921 0.904 0.934 0.488 1.35 1.6
Beta *Weight 0.049243243 0.099567568 0.146594595 0.0252432 0.131891892 0.364864865 0.172972973 0.990378378

So, Portfolio Beta =.99 (Ans)

Ans(4). Total amount for investment is $10,000,000,

Let Amount invested in YHOO is x. So, Amount invested in MGM is (10000000-x)

Required Rate of Return is 50%

Therefor ,Reqired Equation is

x*28.58% +(10000000-x)*80%=10000000*50%

0.2858x-.80x=5000000-8000000

0.5142x=3000000

So, x= 3000000/.5142= $5834305

Amount invested in YHOO =$5834305 Ans.

And Amount invested in MGM=(10000000-x)=(10000000-5834305)=$4165695 Ans.